# Multi-agent Embodied AI: Advances and Future Directions

### v1

@@ 5 The Challenges and Future Works
Building upon the theoretical foundations provided by Markov games, MARL has introduced various frameworks for effectively modeling cooperation among agents in complex environments [245]. Methods such as value decomposition (e.g., VDN [170], QMIX [145]) have improved scalability by decomposing joint objectives into individual utilities, while counterfactual reasoning approaches like COMA [49] have significantly enhanced attribution accuracy for individual agent contributions. Advances in transfer learning and networked MARL have mitigated challenges posed by agent heterogeneity and environmental nonstationarity. Additionally, the integration of control theory and game theory has fostered effective coordination strategies, exemplified by consensus protocols and distributed task allocation methods [33]. However, embodied multi-agent AI introduces unique theoretical challenges, including asynchronous sensing, delayed actions, limited observability, communication constraints, and substantial heterogeneity, complicating both theoretical formulation and practical implementation. Furthermore, while large generative models, particularly LLMs, offer significant potential in facilitating planning and inter-agent communication, their theoretical characteristics—such as stability, generalization, and interpretability—remain inadequately understood [167]. To address these limitations, future research should explore new theoretical paradigms specifically tailored to embodied MAS, leveraging approaches such as causal inference to uncover inter-agent dependencies, complex systems theory to understand emergent behaviors, and hierarchical, bio-inspired coordination frameworks adaptable to real-world complexity.
@@
Recent developments have solidified MARL’s role as a foundational framework for cooperative and competitive multi-agent scenarios. Notable algorithms, including QMIX [145] with its monotonic value decomposition and MAPPO [223], an adaptation of proximal policy optimization, have demonstrated compelling results in benchmark environments such as SMAC and Google research football (GRF)[232]. Nevertheless, these successes predominantly rely on the CTDE paradigm[4], assuming ideal conditions and unrestricted access to agent actions—conditions rarely achievable in embodied scenarios. Physical deployment introduces significant complexities such as sensor noise, limited actuation, delayed feedback, and partial observability. Moreover, existing multi-agent learning frameworks typically struggle with generalization beyond training distributions and have limited scalability when dealing with heterogeneous teams. Embodied multi-agent tasks involve multi-modal sensory data, diverse capabilities, and dynamic interactions, necessitating alternative algorithmic frameworks beyond traditional paradigms such as CTDE, DTDE (decentralized training and execution), or CTCE (centralized training and execution). Promising future directions include hierarchical coordination structures, agent-grouping mechanisms, and the integration of structured priors or classic control theories to achieve scalable and robust performance in embodied MAS.
@@
Most multi-agent learning methodologies have been developed in simulation or gaming environments, benefiting from low-cost, highly repeatable scenarios that allow extensive policy optimization through frequent interaction sampling. However, embodied multi-agent tasks exponentially increase complexity, severely reducing sample efficiency and complicating joint exploration due to vast state and policy spaces [3]. These challenges are amplified in real-world contexts, where each interaction incurs considerable time, financial cost, and hardware wear [229]. Although recent advancements in single-agent embodied learning—such as extracting knowledge from large multimodal models like LLMs [23], employing world-model-based simulations for policy rollouts [114], and leveraging offline data-driven policy calibration [84]—have improved sample efficiency, these strategies often struggle in multi-agent contexts owing to complex interactions and non-stationary dynamics. To overcome these barriers, specialized multi-agent world models accurately simulating interactive dynamics, exploration strategies informed by structured prior knowledge, rapid initialization through IL, and meta-learning for improved generalization are necessary. Additionally, developing robust methodologies supporting effective multi-task and sim-to-real transfer is vital for real-world applicability.
@@
Recent advances in large-scale pretrained models, such as GPT-4, PaLM, CLIP, SAM, and Gemini, have significantly reshaped language, vision, and multimodal learning landscapes by offering powerful capabilities in representation, perception, and reasoning [75]. These foundation models, trained on vast and diverse datasets, provide rich prior knowledge and robust cross-modal alignment, presenting promising avenues for enhancing embodied AI. Leveraging these models allows agents to attain deeper semantic understanding, superior generalization, and more adaptive interactions. However, directly deploying these foundation models in embodied multi-agent scenarios remains challenging [30], as they typically derive from static, single-agent contexts and lack inductive biases essential for dynamic interactions involving asynchronous communication, partial observability, tightly coupled policies, and non-stationarity. Furthermore, embodied MAS usually involve high-dimensional inputs, extensive action spaces, and sparse feedback, limiting adaptability and efficiency. Insights from recent single-agent embodied learning—such as few-shot adaptation, multimodal joint pretraining, and simulation-driven data augmentation—provide valuable transferable knowledge [115]. Future research should focus on developing scalable multi-agent pretraining paradigms, integrating foundation models with reinforcement learning and graph-based coordination methods, and enhancing sim-to-real transfer, thereby forming robust and generalizable frameworks for multi-agent embodied AI in complex, open-ended environments.
@@
Unlike stable and closed settings commonly assumed in embodied AI, open-environment scenarios [258] present significant challenges for multi-agent embodied systems due to their inherently dynamic, uncertain, and non-stationary characteristics. Agents operating within such environments often encounter unpredictable sensory inputs, shifting reward structures, evolving tasks, and frequently changing agent populations. For instance, scenarios may include noisy observations, delayed or disrupted feedback, and task transitions ranging from navigation and exploration to cooperative or competitive interactions. Additionally, team compositions are rarely fixed, with teammates and adversaries changing unpredictably, further complicating coordination. Human involvement introduces extra layers of uncertainty due to diverse, less predictable behaviors. Such variability undermines assumptions about consistency between training and deployment environments, revealing the fragility of policies optimized under stable conditions [232]. Thus, achieving effective performance in open environments requires agents with robust policies resilient to sensory and environmental disturbances, continual learning capabilities that integrate new experiences without compromising previously acquired skills, and contextual reasoning abilities for dynamic interaction interpretation and coordination with novel agents. Methods enabling rapid strategy adaptation, such as policy reuse, latent variable modeling, and meta-learning, become indispensable for generalizing quickly to previously unseen circumstances. Although recent advances in robust MARL, lifelong learning, and context-aware decision-making are promising, critical challenges remain unsolved. Future research should prioritize accurate modeling of open-environment dynamics, anticipating heterogeneous agent and human behaviors, and enabling real-time algorithmic adaptation.
@@
Establishing rigorous and comprehensive evaluation frameworks is crucial yet challenging for advancing multi-agent embodied AI. Recent platforms, including CHAIC [39] (further benchmarks in Table 4), have provided valuable resources but often suffer from limited representational complexity and insufficient ecological validity. Current benchmarks predominantly focus on single-modality tasks, neglecting essential multi-modal signals (vision, language, audio, video) necessary for realistic perception and decision-making. Furthermore, benchmarks commonly involve small-scale, homogeneous MAS, inadequately representing the complexity encountered in heterogeneous MAS comprising diverse embodiments such as drones, robotic arms, autonomous vehicles, and quadrupedal robots. Additionally, idealized assumptions—such as ignoring partial observability, communication latency, adversarial behaviors, and asynchronous execution—widen the sim-to-real gap, undermining policy robustness and transferability. To address these limitations, future research should aim to develop general-purpose, modular, and scalable evaluation frameworks analogous to Gym, MuJoCo, or PyMARL [151], incorporating multi-modal interactions, heterogeneous coordination, and reproducibility features. Physical testbeds, including robot soccer platforms (e.g., RoboCup [44]) and drone swarm arenas, are also essential to validate real-world feasibility and bridge sim-to-real discrepancies. Moreover, the absence of standardized and interpretable evaluation metrics remains unresolved, as existing measures (e.g., coordination scores, network-based metrics) often lack consistency and generalizability. Therefore, developing unified benchmarks with clearly defined tasks, comprehensive evaluation criteria (robustness, scalability, energy efficiency, behavioral diversity), open-access leaderboards, and formal verification methods is imperative, especially for safety-critical applications.
@@
Multi-agent embodied AI holds immense promise across diverse domains, including robotics, education, healthcare, military operations, interactive simulations, and smart-city infrastructures. In robotics, particularly within industrial manufacturing, warehouse logistics, and autonomous driving, successful deployment necessitates robust real-time coordination, reliable collision avoidance, and adaptable platforms, emphasizing algorithmic generalizability. Educational applications involving virtual agent teams demand controlled, socially appropriate behaviors, thus placing substantial requirements on advanced natural language processing and affective computing capabilities. In healthcare, integrating robotic agents within medical teams highlights the necessity for reliable decision-making, robust privacy safeguards, and strict adherence to ethical standards. Military applications, such as coordinated drone swarms, further underline these challenges, stressing policy robustness under adversarial conditions and the crucial role of human oversight to mitigate ethical and operational risks associated with autonomy. Interactive simulations and competitive gaming environments, exemplified by platforms like OpenAI Five [9] and AlphaStar [186], have provided valuable benchmarks but face limitations due to high computational requirements and difficulties in transferring learned strategies to real-world contexts. Smart-city scenarios, including traffic management and energy-grid optimization, involve extensive interactions among numerous autonomous agents, necessitating highly reliable, efficient, and safe solutions with direct societal impacts. Addressing these diverse yet interconnected challenges requires integrating expert domain knowledge, rigorous validation procedures, and comprehensive safety frameworks, enabling practical real-world deployments of multi-agent embodied AI systems.
@@
While the previously discussed sections highlight significant challenges within multi-agent embodied AI, several additional fundamental issues warrant deeper investigation. First, multi-modal perception and collaborative learning remain underexplored, with current research typically focusing on isolated sensory modalities and rarely effectively integrating vision, audio, and language inputs. Future studies should thus prioritize developing robust cross-modal fusion methods and efficient communication protocols to handle modality disparities, information inconsistencies, and latency issues [41]. Second, social learning presents a promising pathway toward emergent collective intelligence. Although approaches such as IL and policy distillation enable agent populations to exhibit coordination surpassing individual capabilities, theoretical foundations and conditions necessary for stable emergent behaviors remain inadequately understood [128]. Lastly, as MAS increasingly gain autonomy, addressing safety and ethical considerations becomes critically important, especially within sensitive or high-stakes environments [150]. Explicit safety constraints, clear ethical guidelines, and transparent explainability mechanisms must be developed to ensure fair, accountable, and interpretable decision-making, thereby minimizing unintended risks arising from complex agent interactions and ensuring trustworthy AI deployments.

### v2

@@ 5 The Challenges and Future Works
Building upon the theoretical foundations provided by Markov games, MARL has introduced various frameworks for effectively modeling cooperative relationships among agents in complex environments [252]. Methods such as value decomposition (e.g., VDN [176], QMIX [150]) have improved scalability by decomposing joint objectives into individual utilities, while counterfactual reasoning approaches like COMA [50] have significantly enhanced attribution accuracy for individual agent contributions. Advances in transfer learning and networked MARL have mitigated challenges posed by agent heterogeneity and environmental non-stationarity. Additionally, the integration of control and game theory has fostered effective coordination strategies, exemplified by consensus protocols and distributed task allocation methods [34]. However, embodied multi-agent intelligence introduces unique theoretical challenges, including asynchronous sensing, delayed actions, limited observability, communication constraints, and substantial heterogeneity, complicating both theoretical formulation and practical implementation. Furthermore, while large generative models, particularly LLMs, offer significant potential in facilitating planning and inter-agent communication, their theoretical characteristics such as stability, generalization and interpretability still remain inadequately understood [173]. To address these limitations, future research should explore new theoretical paradigms specifically tailored to embodied MAS, leveraging approaches such as causal inference to uncover inter-agent dependencies, complex systems theory to understand emergent behaviors, and hierarchical, bio-inspired coordination frameworks adaptable to real-world complexity.
@@
Recent developments have solidified the role of MARL as a foundational framework for cooperative and competitive multi-agent scenarios. Notable algorithms, including QMIX [150] with its monotonic value decomposition and MAPPO [230], an adaptation of proximal policy optimization, have demonstrated compelling results in benchmark environments such as SMAC and Google research football (GRF)[239]. Nevertheless, these successes predominantly rely on the CTDE paradigm[4], assuming ideal conditions and unrestricted access to agent actions—conditions rarely achievable in embodied scenarios. Physical deployment introduces significant complexities such as sensor noise, limited actuation, delayed feedback, and partial observability. Moreover, existing multi-agent learning frameworks typically struggle with generalization beyond training distributions and have limited scalability when dealing with heterogeneous teams. Embodied multi-agent tasks involve multi-modal sensory data, diverse capabilities, and dynamic interactions, necessitating alternative algorithmic frameworks beyond traditional paradigms such as CTDE, DTDE (decentralized training and execution), or CTCE (centralized training and execution). Promising future directions include hierarchical coordination structures, agent-grouping mechanisms, and the integration of structured priors or classic control theories to achieve scalable and robust performance in embodied MAS.
@@
Most multi-agent learning methodologies have been developed in simulation or gaming environments, benefiting from low-cost, highly repeatable scenarios that allow extensive policy optimization through frequent interaction sampling. However, embodied multi-agent tasks exponentially increase complexity, severely reducing sample efficiency and complicating joint exploration due to vast state and policy spaces [3]. These challenges are amplified in real-world contexts, where each interaction incurs considerable time, financial cost, and hardware wear [236]. Although recent advancements in single-agent embodied learning—such as extracting knowledge from large multi-modal models like LLMs [24], employing world-model-based simulations for policy rollouts [117], and leveraging offline data-driven policy calibration [85]—have improved sample efficiency, these strategies often struggle in multi-agent contexts owing to complex interactions and non-stationary dynamics. To overcome these barriers, specialized multi-agent world models accurately simulating interactive dynamics, exploration strategies informed by structured prior knowledge, rapid initialization through IL, and meta-learning for improved generalization are necessary. Additionally, developing robust methodologies supporting effective multi-task and sim-to-real transfer is vital for real-world applicability.
@@
Recent advances in pre-trained large models, such as GPT-4, PaLM, CLIP, SAM and Gemini, have significantly reshaped language, vision and multi-modal learning landscapes by offering powerful capabilities in representation, perception, and reasoning [76]. These foundation models, trained on vast and diverse datasets, provide rich prior knowledge and robust cross-modal alignment, presenting promising avenues for enhancing embodied AI. Leveraging these models allows agents to attain deeper semantic understanding, superior generalization, and more adaptive interactions. However, directly deploying these foundation models in embodied multi-agent scenarios remains challenging [31], as they typically derive from static, single-agent contexts and lack inductive biases essential for dynamic interactions involving asynchronous communication, partial observability, tightly coupled policies, and non-stationarity. Furthermore, embodied MAS usually involve high-dimensional multi-modal inputs, extensive action spaces and sparse feedback, limiting the adaptability and efficiency of the learning algorithms. Insights from recent single-agent embodied learning—such as few-shot adaptation, multi-modal joint pretraining and simulation-driven data augmentation provide valuable transferable knowledge [118] for multi-agent research. Future research should focus on developing scalable multi-agent pretraining paradigms, integrating foundation models with reinforcement learning and graph-based coordination methods, and enhancing sim-to-real transfer, thereby forming robust and generalizable frameworks for multi-agent embodied AI in complex, open-ended environments.
@@
Unlike the stable and closed task settings that commonly assumed in embodied AI research, scenarios with open environments [267] present significant challenges for embodied MAS due to their inherently dynamic, uncertain and non-stationary characteristics. Agents operating within such environments often encounter unpredictable sensory inputs, shifting reward structures, evolving tasks and frequently changing agent populations. For instance, the environments may include noisy observations, delayed or disrupted feedback and task transitions ranging from navigation and exploration to cooperative or competitive interactions. Additionally, the shifting team compositions with teammates and adversaries changing unpredictably further complicate coordinations. The human involvement also introduces extra layers of uncertainty due to diverse, less predictable behaviors. Such variability undermines the consistency between training and deployment environments, revealing the fragility of policies [239]. Thus, achieving effective performance in open environments requires agents with robust policies resilient to sensory and environmental disturbances, continual learning capabilities that integrate new experiences without compromising previously acquired skills and contextual reasoning abilities for dynamic interactions and coordinations with novel agents. Methods enabling rapid strategy adaptation such as policy reuse, latent variable modeling and meta-learning have become indispensable for generalizing quickly to unseen circumstances. Although recent advances in robust MARL, lifelong learning, and context-aware decision-making are promising, critical challenges remain unsolved. Future research should prioritize accurate modeling of open-environment dynamics, anticipating heterogeneous agent and human behaviors, and enabling real-time algorithmic adaptation.
@@
Establishing rigorous and comprehensive evaluation frameworks is crucial yet challenging for advancing multi-agent embodied AI. Recent platforms, including CHAIC [40] (further benchmarks in Table 4), have provided valuable resources but often suffer from limited representational complexity and insufficient ecological validity. Current benchmarks predominantly focus on single-modality tasks, neglecting essential multi-modal signals (vision, language, audio, video) necessary for realistic perception and decision-making. Furthermore, benchmarks commonly involve small-scale, homogeneous MAS, inadequately representing the complexity encountered in heterogeneous MAS comprising diverse embodiments such as drones, robotic arms, autonomous vehicles, and quadrupedal robots. Additionally, idealized assumptions such as ignoring partial observability, communication latency, adversarial behaviors and asynchronous execution widen the sim-to-real gap, undermining policy robustness and transferability. Subsequently, future research should aim to develop general-purpose, modular and scalable evaluation frameworks analogous to Gym, MuJoCo or PyMARL [157], incorporating multi-modal interactions, heterogeneous coordination and reproducibility features. Physical testbeds, including robot soccer platforms (e.g., RoboCup [45]) and drone swarm arenas are also essential to validate real-world feasibility and bridge sim-to-real gaps. Moreover, the absence of standardized and interpretable evaluation metrics remains unresolved, as existing measures (e.g., coordination scores, network-based metrics) often lack consistency and generalizability. Therefore, developing unified benchmarks with clearly defined tasks, comprehensive evaluation criteria (robustness, scalability, energy efficiency, behavioral diversity), open-access leader-boards, and formal verification methods is imperative, especially for safety-critical applications.
@@
Multi-agent embodied AI holds immense promise across diverse domains including robotics, education, healthcare, military operations, interactive simulations and smart-city infrastructures. In robotics, particularly within industrial manufacturing, warehouse logistics, and autonomous driving, successful deployment necessitates robust real-time coordination, reliable collision avoidance, and adaptable platforms, emphasizing algorithmic generalizability. Educational applications involving virtual agent teams demand controlled, socially appropriate behaviors, thus placing substantial requirements on advanced natural language processing and affective computing capabilities. In healthcare, integrating robotic agents within medical teams highlights the necessity for reliable decision-making, robust privacy safeguards, and strict adherence to ethical standards. Military applications, such as coordinated drone swarms, further underline these challenges, stressing policy robustness under adversarial conditions and the crucial role of human oversight to mitigate ethical and operational risks associated with autonomy. Interactive simulations and competitive gaming environments, exemplified by platforms like OpenAI Five [9] and AlphaStar [192], have provided valuable benchmarks but face limitations due to high computational requirements and difficulties in transferring learned strategies to real-world contexts. Smart-city scenarios, including traffic management and energy-grid optimization, involve extensive interactions among numerous autonomous agents, necessitating highly reliable, efficient, and safe solutions with direct societal impacts. Addressing these diverse yet interconnected challenges requires integrating expert domain knowledge, rigorous validation procedures, and comprehensive safety frameworks, enabling practical real-world deployments of embodied intelligent MAS.
@@
While the previously sections highlight significant challenges within multi-agent embodied AI, several additional fundamental issues warrant deeper investigation. First, multi-modal perception and collaborative learning remain underexplored, with current research typically focusing on isolated sensory modalities and rarely effectively integrating vision, audio and language inputs. Future studies should thus prioritize developing robust cross-modal fusion methods and efficient communication protocols to handle modality disparities, information inconsistencies and latency issues [42]. Second, social learning presents a promising pathway toward emergent collective AI. Although approaches such as IL and policy distillation enable agent populations to exhibit coordination surpassing individual capabilities, theoretical foundations and conditions necessary for stable emergent behaviors remain inadequately understood [132]. Lastly, as MAS are increasingly gain autonomy, addressing safety and ethical considerations becomes critically important, especially within sensitive or high-stakes environments [156]. Explicit safety constraints, clear ethical guidelines and transparent explainability mechanisms must be developed to ensure fair, accountable and interpretable decision-making, thereby minimizing unintended risks arising from complex agent interactions and ensuring trustworthy AI deployments.